# Coherence-Mediated Quantum Thermometry in a Hybrid Circuit-QED Architecture

### v1

@@ Abstract
Quantum thermometry plays a critical role in the development of low-temperature sensors and quantum information platforms. In this work, we propose and theoretically analyze a hybrid circuit quantum electrodynamics architecture in which a superconducting qubit is dispersively coupled to two distinct bosonic modes: one initialized in a weak coherent state and the other coupled to a thermal environment. We show that the qubit serves as a sensitive readout of the probe mode, mapping the interference between thermal and coherent photon-number fluctuations onto measurable dephasing. This mechanism enables enhanced sensitivity to sub-millikelvin thermal energy fluctuations through Ramsey interferometry. We derive analytic expressions for the qubit coherence envelope, compute the quantum Fisher information for temperature estimation, and demonstrate numerically that the presence of a coherent reference amplifies the qubit’s sensitivity to small changes in thermal photon occupancy. Our results establish a new paradigm for quantum-enhanced thermometry and provide a scalable platform for future calorimetric sensing in high-energy physics and quantum metrology.
@@ I Introduction
Quantum thermometry has emerged as a critical capability in modern quantum science and technology. It plays a central role in the calibration of cryogenic environments, the characterization of quantum devices and sensors, and the detection of weak energy events in low-temperature physics [1, 2, 3, 4]. Beyond practical applications, precise temperature measurement also provides a platform for exploring fundamental thermodynamic limits in the quantum regime. Superconducting quantum circuits, particularly those operating in the microwave domain, offer a versatile and scalable platform for implementing such thermometric protocols [5, 6]. In these systems, qubits serve as exquisitely sensitive probes of their electromagnetic environment and can detect thermal photon populations through their decoherence dynamics [7, 8, 9, 10, 11].
@@
The most common approach to qubit-based thermometry exploits the dispersive interaction between a qubit and a thermalized cavity mode. In the dispersive regime, photon-number fluctuations induce pure dephasing of the qubit, which can be monitored via Ramsey or spin-echo sequences [12, 13]. The decay of qubit coherence directly encodes the thermal occupancy $\bar{n}$ of the mode and thereby its effective temperature. This single-mode dephasing protocol has been demonstrated in a variety of circuit quantum electrodynamics (cQED) architectures [14, 15, 16] and is particularly effective when $\bar{n}\gtrsim 1$. However, in the ultra-low-temperature limit $\bar{n}\ll 1$, corresponding to mode temperatures below $50$ mK at GHz frequencies, the variance of Bose–Einstein photon statistics becomes exponentially suppressed, and the sensitivity of this method is fundamentally limited by the vanishing signal-to-noise ratio of thermal fluctuations.
@@
Here we introduce a two-mode quantum thermometry protocol that overcomes this low-temperature sensitivity floor by leveraging coherent–thermal interference. Specifically, we consider a hybrid cQED architecture in which a superconducting qubit is dispersively coupled to two bosonic modes: a high-$Q$ 3D cavity initialized in a weak coherent state, and a low-$Q$ planar resonator coupled to a thermal environment. The coherent mode acts as a stable phase reference, while the thermal mode functions as the sensing arm. The qubit, coupled to the probe cavity, serves as a phase-sensitive readout that acquires a stochastic phase shift determined by the joint photon-number fluctuations of both modes. Therefore, the qubit’s coherence decay directly reflects the interference between their photon statistics.
@@
This architecture realizes a novel interferometric thermometry scheme in which small thermal signals are amplified by embedding them in the phase dynamics of a coherent reference. We analytically derive the coherence envelope of the probe mode as a function of interaction time, thermal photon number, and coherent amplitude, and show how it is faithfully mapped onto qubit Ramsey dephasing through dispersive readout. The resulting dephasing dynamics depend not only on the variances of each mode individually but also on their interference, enabling thermal signal amplification through coherent–thermal correlations. To quantify the resulting sensitivity, we calculate the quantum Fisher information (QFI) associated with qubit coherence measurements and identify optimal working points that maximize temperature resolution [17, 18, 19, 20].
@@
Numerical simulations based on exact unitary interaction validate the analytic model and show that sub-millikelvin sensitivity is, in principle, achievable with realistic circuit parameters. The proposed architecture is compatible with existing cQED platforms, including hybrid designs that combine planar resonators and qubits with 3D cavities [21, 22, 23], and does not require photon counting or non-Gaussian state preparation. These features make it a scalable and noise-resilient foundation for thermometry at ultra-low energies. Potential applications range from on-chip quantum diagnostics [24] and cryogenic calorimetry [25] to thermal signal amplification in rare-event detection platforms such as axion or hidden-photon dark matter searches [26, 27]. Taken together, these results establish coherent–thermal interferometry as a promising route toward quantum-limited thermometric sensing in superconducting circuits.
@@ II Theory: Thermal-to-Coherent Mapping
Quantum thermometry aims to extract information about temperature from quantum systems with minimal disturbance and maximal sensitivity. In this work, we consider a hybrid bosonic system where a thermal field is entangled with a coherent reference field. This interaction, acting as a noise-to-phase transduction mechanism, enables temperature-dependent fluctuations in the thermal field to be imprinted onto a well-controlled coherent mode, thereby amplifying otherwise weak thermal signals and allowing them to be efficiently detected by monitoring the qubit’s coherence. Below we present a detailed theoretical framework for this thermal-to-coherent mapping mechanism.
@@ II.1 System Model
Figure 1: Schematic of the proposed thermometry architecture. A low-$Q$ thermal resonator $\hat{a}$ is thermalized by a bath at temperature $T$. Through an engineered cross-Kerr interaction $\lambda$, realized with two fixed-frequency transmons ($Q_{1}$, $Q_{2}$), fluctuations of $\hat{a}$ are mapped onto a high-$Q$ 3D cavity probe mode $\hat{b}$. A sensing transmon couples strongly to the probe with dispersive strength $\chi_{b}$ for readout, while any residual coupling $\chi_{a}$ to the thermal mode acts as a parasitic dephasing channel. Temperature information encoded in the probe can be extracted either via qubit Ramsey (coherence-mediated) or direct heterodyne detection (phase-shift).
As shown in Fig. 1, we consider a hybrid system comprising a qubit coupled to two bosonic modes:
Mode $\hat{a}$: a thermal field at temperature $T$, initialized in a thermal state $\rho_{\mathrm{th}}(T)$ with mean photon number $\bar{n}_{a}$. A low-quality 2D resonator, such as a CPW or lumped-element LC resonator, can be used to realize this mode.
Mode $\hat{b}$: a coherent reference field prepared in a coherent state $|\alpha\rangle$ with amplitude $\alpha$. This mode can be implemented using an ultra-high-quality 3D cavity.
The qubit interacts dispersively with both modes. In addition, a nonlinear cross-Kerr interaction mediates a joint coupling between the thermal and coherent modes. The total interaction Hamiltonian is given by
@@ II.2 Lamb-Like Shift: Thermal-to-Coherent Mapping
The cross-Kerr interaction, $\lambda\hat{n}_{a}\hat{n}_{b}$, couples the thermal mode $\hat{a}$ to the coherent probe mode $\hat{b}$, such that thermal photon-number fluctuations induce frequency shifts in the coherent mode [31, 32]. While the traditional Lamb shift arises as a second-order correction to a quantum system’s energy levels due to vacuum fluctuations, the present scenario exhibits a reversed structure: it is the coherent field, rather than the qubit, that experiences environment-induced shifts. Here, the thermal mode acts as an effective fluctuating environment for the probe.
@@
where the thermal occupation $\hat{n}_{a}$ follows Bose–Einstein statistics: $\bar{n}_{a}(T)=1/[\text{exp}(\hbar\omega_{a}/k_{B}T)-1].$ Due to the probabilistic nature of thermal fluctuations, each realization of $n_{a}$ photons shifts the frequency of the coherent mode by $\lambda n_{a}$, resulting in a distribution of frequency shifts centered at $\omega_{b}+\lambda\bar{n}_{a}$.
This random frequency shift leads to stochastic phase accumulation. For a single realization with $n_{a}$ photons, the phase shift after time $\tau$ is deterministic: $\phi_{b}(\tau,T)=\lambda n_{a}\tau$. Across the thermal ensemble, however, this becomes a random variable: $\phi_{b}=\int_{0}^{\tau}\delta\omega_{b}(t)\,dt,$ where $\delta\omega_{b}(t)$ captures the instantaneous fluctuation in frequency due to the thermal photon number. This process defines a pure dephasing channel for the coherent state: it causes a gradual loss of phase coherence without energy dissipation.
@@
II.3 Strategy 1: Coherence-mediated envelope and QFI
After interacting with the thermal mode via a cross-Kerr coupling, the coherent probe acquires temperature-dependent phase noise. This dephased coherent state is interrogated by a dispersively coupled qubit using a Ramsey sequence [33, 34]. The qubit’s coherence reflects the fluctuating probe phase:
$|\psi_{q}(\tau)\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+e^{i\phi_{b}}|1\rangle\right),$ (3)
where the accumulated phase $\phi_{b}=\lambda n_{a}\tau$ originates from thermal fluctuations and is therefore temperature dependent.
We model the probe mode as undergoing random phase kicks induced by photon-number fluctuations in the thermal mode. Since $\phi_{b}\propto n_{a}$, its statistics are inherited from the Bose–Einstein distribution of $n_{a}$, with mean $\bar{n}_{a}$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}(\bar{n}_{a}+1)$. Accordingly, the probe phase has mean $\langle\phi_{b}\rangle=\lambda\tau\bar{n}_{a}$ and variance
@@ II.3 Strategy 1: Coherence-mediated envelope and QFI
$\rho_{b}(T)=\int d\phi\,\frac{e^{-\phi^{2}/2\sigma_{\phi}^{2}}}{\sqrt{2\pi\sigma_{\phi}^{2}}}\,|\alpha e^{i\phi}\rangle\langle\alpha e^{i\phi}|,$ (5)
@@
Figure 2: Simulated coherence-based QFI $\mathcal{F}_{C}$ and corresponding temperature sensitivity $\delta T$, with thermal frequency $\omega_{a}/2\pi=1~\mathrm{GHz}$, coherent amplitude $\alpha=2.0$, and cross-Kerr coupling $\lambda/2\pi=50~\mathrm{kHz}$. (a) Heatmap of $\mathcal{F}_{C}$ as a function of thermal bath temperature $T$ and interaction time $\tau$. The red dashed line indicates the optimal $\tau(T)$ that maximizes $\mathcal{F}_{C}$ for each temperature. (b) Minimum detectable temperature change $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}}$ (blue solid line) as a function of temperature, computed at the optimal $\tau$ (green dashed line). A sensitivity of approximately $60~\mu\mathrm{K}$ is achieved near $T\sim 10~\mathrm{mK}$ with $\tau=10~\mu\mathrm{s}$. In this simulation, we set the number of measurement repetitions to $\nu=10^{4}$.
@@
$\mathcal{V}_{C}(\tau,T)\equiv C(\tau,T)=\exp\left[-2|\alpha|^{2}\left(1-e^{-\Gamma_{\phi}}\right)\right],$ (6)
with the effective dephasing rate $\Gamma_{\phi}=\sigma_{\phi}^{2}\approx(\lambda\tau)^{2}\bar{n}_{a}$ (Appendix B).
@@
$\mathcal{F}_{C}(T)=\frac{\left[2\alpha^{2}e^{-\Gamma_{\phi}}(\lambda\tau)^{2}\,\partial_{T}\bar{n}_{a}\right]^{2}C^{2}}{1-C^{2}},$ (7)
where $\partial_{T}\bar{n}_{a}=\hbar\omega_{a}/\!\left[\bar{n}_{a}(\bar{n}_{a}+1)k_{B}T^{2}\right]$. Eq. 7 establishes the fundamental precision bound of the thermometer [17, 19, 20, 18, 39].
Physically, the probe converts temperature-dependent photon-number fluctuations into measurable phase diffusion. In this coherence-mediated model, the QFI increases with $\alpha^{2}$ but is simultaneously limited by visibility decay in the denominator. In the weak-dephasing regime ($\Gamma_{\phi}\ll 1$), one finds $\mathcal{F}_{C}\;\approx\;\alpha^{2}(\lambda\tau)^{2}\,\bar{n}_{a}\left(\hbar\omega_{a}/k_{B}T^{2}\right)^{2}.$ Thus, the coherent-probe scheme provides a tunable transduction “gain” via $\alpha$ (and $\tau$), but this gain is bounded in practice by visibility decay, available probe power, bandwidth, and backaction. We therefore interpret the improvement as a resource-assisted sensitivity, which is achieved by allocating probe photons to strengthen the temperature-to-phase conversion, rather than as an unlimited amplification.
Figure 2 shows a simulated heatmap of $\mathcal{F}_{C}$, highlighting the optimal interaction time $\tau(T)$ shown as dashed line in panel(a). Temperature sensitivity is defined as $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}(T)},$ with $\nu$ the number of independent repetitions. For realistic parameters ($\lambda/2\pi=50~\mathrm{kHz}$, $\omega_{a}/2\pi=1~\mathrm{GHz}$, and $\alpha=2$), the thermometer achieves a sensitivity $\delta T\sim 60~\mu\mathrm{K}$ near $T=10~\mathrm{mK}$, with an optimal interaction time around $\tau\sim 10~\mu\mathrm{s}$ and measurement repetition $\nu=10^{4}$. Such timescales are readily supported by high-Q 3D cavities, which exhibit photon lifetimes up to tens of milliseconds [40, 41, 23].
II.4 Strategy 2: Probe Phase-Shift Tracking

### v2

@@ Abstract
Quantum thermometry plays a critical role in the development of low-temperature sensors and quantum information platforms. In this work, we propose and analyze a hybrid circuit quantum electrodynamics architecture in which a superconducting qubit is dispersively coupled to two distinct bosonic modes: one initialized in a weak coherent state as a phase reference and information buffer and the other coupled to a thermal environment. We show that the qubit serves as a sensitive readout of the probe mode, mapping the interplay between thermal and coherent photon-number fluctuations onto measurable dephasing. This coherence-mediated mechanism enables improved sensitivity to thermal energy fluctuations in the sub-millikelvin regime through Ramsey interferometry. We derive analytic expressions for the probe coherence envelope, compute the quantum Fisher information for temperature estimation, and demonstrate numerically that the presence of a coherent reference enhances the qubit’s sensitivity to small changes in thermal photon occupancy. Our results establish a coherence-enabled approach to thermometry and provide a scalable platform for future calorimetric sensing in high-energy physics and quantum metrology.
@@ I Introduction
Quantum thermometry has emerged as a critical capability in modern quantum science and technology. It plays a central role in the calibration of cryogenic environments, the characterization of quantum devices and sensors, and the detection of weak energy events in low-temperature physics [11, 16, 59, 45]. Beyond practical applications, precise temperature measurement also provides a platform for exploring fundamental thermodynamic limits in the quantum regime. Superconducting quantum circuits, particularly those operating in the microwave domain, offer a versatile and scalable platform for implementing such thermometric protocols [19, 38]. In these systems, qubits serve as exquisitely sensitive probes of their electromagnetic environment and can detect thermal photon populations through their decoherence dynamics [34, 55, 60, 58].
@@
The most common approach to qubit-based thermometry exploits the dispersive interaction between a qubit and a thermalized cavity mode. In the dispersive regime, photon-number fluctuations induce pure dephasing of the qubit, which can be monitored via Ramsey sequences [57, 8]. The decay of qubit coherence directly encodes the thermal occupancy $\bar{n}_{th}$ of the mode and thereby its effective temperature. This single-mode dephasing protocol has been demonstrated in a variety of circuit quantum electrodynamics (cQED) architectures [41, 46, 7, 64, 44, 42] and is particularly effective when $\bar{n}_{th}\gtrsim 1$. However, in the ultra-low-temperature limit $\bar{n}_{th}\ll 1$, corresponding to mode temperatures below $50$ mK at GHz frequencies, the variance of Bose–Einstein photon statistics becomes exponentially suppressed, and the sensitivity of this method is fundamentally limited by the vanishing signal-to-noise ratio of thermal fluctuations. Beyond qubit-based approaches, a range of established cryogenic thermometry techniques, such as Coulomb-blockade thermometers and proximity-based superconducting thermometers, provide robust calibration standards and performance across complementary temperature ranges [52, 63, 31].
@@
Here, we introduce the coherence-mediated quantum thermometry (CMQT), a two-mode quantum thermometry protocol that overcomes this low-temperature sensitivity floor by leveraging coherent-thermal interference. We consider a hybrid cQED architecture in which a superconducting qubit is dispersively coupled to two bosonic modes: a high-$Q$ 3D cavity initialized in a weak coherent state, and a low-$Q$ planar resonator coupled to a thermal environment. The coherent mode acts as a stable phase reference, while the thermal absorber mode functions as the sensing arm. The qubit, coupled to the probe cavity, serves as a phase-sensitive readout whose Ramsey phase acquires a stochastic shift determined by the joint photon-number fluctuations of both modes. As a result, the qubit’s coherence decay directly reflects the interference between their photon statistics. In this framing, the key capability of CMQT is the separation of roles: the absorber is engineered for strong and well-characterized thermalization, the probe acts as a phase-coherent information buffer, and the qubit is only engaged during a short mapping window. This modularity enables independent optimization of the thermal coupling and bandwidth, the storage time and technical noise filtering, and the qubit coherence and readout fidelity, a set of trade-offs that is difficult to realize when the qubit simultaneously serves as sensor, memory, and readout.
@@
This architecture realizes an interferometric thermometry scheme in which small thermal signals are amplified by embedding them in the phase dynamics of a coherent reference. We analytically derive the coherence envelope of the probe mode as a function of interaction time, thermal photon number, and coherent amplitude, and show how it is faithfully mapped onto qubit Ramsey dephasing through dispersive readout. The resulting dephasing dynamics depend not only on the variances of each mode individually but also on their interference, enabling thermal signal amplification through coherent–thermal correlations. To quantify the resulting sensitivity, we calculate the quantum Fisher information (QFI) associated with qubit coherence measurements and identify optimal working points that maximize temperature resolution [46, 17, 40, 67]. This added hybrid complexity is motivated by the ultra-low-occupation regime $\bar{n}_{a}\!\ll\!1$, where conventional single-mode dispersive thermometry exhibits a parametrically small temperature responsivity.
Figure 1: Schematic of the proposed thermometry architecture. A low-$Q$ thermal resonator $\hat{a}$ is equilibrated with a bath at temperature $T$. Its photon-number fluctuations are imprinted to a high-$Q$ 3D cavity probe mode $\hat{b}$ via an engineered cross-Kerr interaction $\lambda$ implemented through two intermediate qubits ($Q_{1}$, $Q_{2}$). A readout transmon qubit couples dispersively to the probe with strength $\chi_{b}$ for phase mapping, while any residual coupling $\chi_{a}$ to the thermal mode introduces parasitic qubit dephasing. The temperature-dependent phase statistics encoded in the probe can be extracted either through a qubit Ramsey measurement or through direct heterodyne detection. Here $\omega$ denotes the mode frequencies, $\kappa$ their linewidths, and $\alpha_{q}$ the qubit anharmonicity.
@@
Numerical simulations based on exact unitary interaction validate the analytic model and show that sub-millikelvin sensitivity is, in principle, achievable with realistic circuit parameters. The proposed architecture is compatible with existing cQED platforms, including hybrid designs that combine planar resonators and qubits with 3D cavities [2, 26, 36], and does not require photon counting or non-Gaussian state preparation. These features make it a scalable and noise-resilient foundation for thermometry at ultra-low energies. Potential applications range from on-chip quantum diagnostics [28] and cryogenic calorimetry [48] to thermal signal amplification in rare-event detection platforms such as axion or hidden-photon dark matter searches [6, 43]. Thus, these results establish coherent–thermal interferometry as a promising route toward quantum-limited thermometric sensing in superconducting circuits. We note practical constraints that bound performance in realistic devices, including maintaining dispersive operation, calibrating AC-Stark/Kerr drifts and drive-induced leakage under coherent drives, characterizing non-thermal absorber noise, and managing long averaging times at ultra-low $T$ with minimal per-shot backaction. Instead, in this work we identify the core CMQT operating point and its regime of advantage using a minimal, transparent model that isolates the essential temperature-to-phase transduction mechanism.
This article is organized as follows. Section II introduces the theoretical model and hybrid circuit-QED architecture, derives the thermal-to-coherent mapping mechanism, and outlines two operating strategies suited to different measurement protocols and temperature ranges. Section III details an experimentally realistic implementation, including mode realization, the engineered cross-Kerr interaction, expected qubit decoherence, and strategies for mitigating excess noise. Section IV benchmarks CMQT against a conventional qubit-only thermometer in which the qubit acts as both sensor and memory, emphasizing the resulting preservation of qubit coherence and the associated sensitivity improvement. Section V summarizes the main results and discusses practical limitations and outlook. Further technical derivations are collected in the Appendices.
@@ II Theory: Thermal-to-Coherent Mapping
Quantum thermometry aims to extract information about temperature from quantum systems with minimal disturbance and maximal sensitivity. In this work, we consider a hybrid bosonic system where a thermal field is coupled with a coherent reference field. This interaction, acting as a noise-to-phase transduction mechanism, enables temperature-dependent fluctuations in the thermal field to be imprinted onto a well-controlled coherent mode, thereby amplifying otherwise weak thermal signals and allowing them to be efficiently detected by monitoring the qubit’s coherence. Below we present a detailed theoretical framework for this thermal-to-coherent mapping mechanism.
@@ II.1 System Model
As shown in Fig. 1, we consider a hybrid system comprising a transmon qubit coupled to two bosonic modes:
Mode $\hat{a}$: a thermal field at temperature $T$, initialized in a thermal state $\rho_{\mathrm{th}}(T)$ with mean photon number $\bar{n}_{a}$. A 2D planar resonator, such as a CPW or lumped-element LC resonator, can be used to realize this mode.
Mode $\hat{b}$: a coherent reference field prepared in a coherent state $|\alpha\rangle$ with amplitude $\alpha$. This mode can be implemented using a high-quality 3D cavity.
A strong dispersive coupling $\chi_{b}$ enables the transmon qubit to map any temperature-dependent phase shift accumulated in the coherent probe mode, whereas the parasitic coupling $\chi_{a}$ to the thermal mode introduces additional dephasing in the qubit.
In addition, a nonlinear cross-Kerr interaction mediates a joint coupling between the thermal and coherent modes. The total interaction Hamiltonian is given by
@@ II.2 Lamb-Like Shift: Thermal-to-Coherent Mapping
The cross-Kerr interaction, $\lambda\hat{n}_{a}\hat{n}_{b}$, couples the thermal absorber $a$ to the coherent probe $b$, such that thermal photon-number fluctuations induce frequency shifts in the coherent mode [4, 54]. While the traditional Lamb shift arises as a second-order correction to a quantum system’s energy levels due to vacuum fluctuations, the present scenario exhibits a reversed structure: it is the coherent field, rather than the qubit, that experiences environment-induced shifts. Here, the thermal mode acts as an effective fluctuating environment for the probe.
@@
where the thermal occupation $\hat{n}_{a}$ follows Bose-Einstein statistics: $\bar{n}_{a}(T)=1/[\text{exp}(\hbar\omega_{a}/k_{B}T)-1].$ Due to the probabilistic nature of thermal fluctuations, each realization of $n_{a}$ photons shifts the frequency of the coherent mode by $\lambda n_{a}$, resulting in a distribution of frequency shifts centered at $\omega_{b}+\lambda\bar{n}_{a}$.
This random frequency shift leads to stochastic phase accumulation. For a single realization with $n_{a}$ photons, the phase shift after the absorber-probe interaction time $\tau$ is deterministic: $\phi_{b}(\tau,T)=\lambda n_{a}\tau$. Across the thermal ensemble, however, this becomes a random variable: $\phi_{b}=\int_{0}^{\tau}\delta\omega_{b}(t)\,dt,$ where $\delta\omega_{b}(t)$ captures the instantaneous fluctuation in frequency due to the thermal photon number. This process defines a pure dephasing channel for the coherent state: it causes a gradual loss of phase coherence without energy dissipation.
@@
II.3 Strategy 1: Coherence-mediated (CM) envelope and QFI
Figure 2: Simulated coherence-based QFI, $\mathcal{F}_{C}$, and corresponding temperature sensitivity, $\delta T$, for a thermal mode with frequency $\omega_{a}/2\pi=1~\mathrm{GHz}$, probe amplitude $\alpha=1.0$, and cross-Kerr coupling $\lambda/2\pi=50~\mathrm{kHz}$. (a) Heatmap of $\mathcal{F}_{C}$ as a function of thermal bath temperature $T$ and interaction time $\tau$. The white dashed curve denotes the optimal interaction time $\tau_{\mathrm{opt}}(T)$ that maximizes $\mathcal{F}_{C}$ at each temperature. (b) Minimum detectable temperature change $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}}$ (red), evaluated along $\tau_{\mathrm{opt}}(T)$ (green). A sensitivity of approximately $\delta T=60~\mu\mathrm{K}$ is achieved at $T=10~\mathrm{mK}$ with $\tau_{\mathrm{opt}}\approx 25~\mu\mathrm{s}$ for $\nu=10^{4}$ repetitions. (c) QFI as a function of interaction time $\tau$ at fixed $T=10~\mathrm{mK}$ for two different cross-Kerr strengths $\lambda$. While the maximum attainable QFI is independent of $\lambda$, the optimal $\tau$ decreases for larger $\lambda$, reflecting faster phase accumulation.
After interacting with the thermal mode via a cross-Kerr coupling, the coherent probe acquires temperature-dependent phase noise. This dephased coherent state is interrogated by a dispersively coupled qubit using a Ramsey sequence [10, 56]. The qubit’s coherence reflects the fluctuating probe phase: $|\psi_{q}(\tau)\rangle=\tfrac{1}{\sqrt{2}}\left(|0\rangle+e^{i\phi_{b}}|1\rangle\right),$ where the accumulated phase $\phi_{b}=\lambda n_{a}\tau$ originates from thermal fluctuations and is therefore temperature dependent.
We model the probe mode as undergoing random phase kicks induced by photon-number fluctuations in the thermal mode under the quasi-static condition ($\kappa_{a}\tau<1$). Since $\phi_{b}\propto n_{a}$, its statistics are inherited from the Bose-Einstein distribution of $n_{a}$, with mean $\bar{n}_{a}$ and variance $\mathrm{Var}(n_{a})=\bar{n}_{a}(\bar{n}_{a}+1)$. Accordingly, the probe phase has mean $\langle\phi_{b}\rangle=\lambda\tau\bar{n}_{a}$ and variance
@@ II.3 Strategy 1: Coherence-mediated (CM) envelope and QFI
$\rho_{b}(T)=\int d\phi\,\tfrac{e^{-\phi^{2}/2\sigma_{\phi}^{2}}}{\sqrt{2\pi\sigma_{\phi}^{2}}}\,|\alpha e^{i\phi}\rangle\langle\alpha e^{i\phi}|,$ (4)
@@
$C(\tau,T)=\exp\left[-2|\alpha|^{2}\left(1-e^{-\Gamma_{\phi}}\right)\right],$ (5)
with the effective dephasing rate $\Gamma_{\phi}=\sigma_{\phi}^{2}/2$ (Appendix B).
@@
$\mathcal{F}_{C}(T)=\frac{\left[\alpha^{2}e^{-\Gamma_{\phi}}(\lambda\tau)^{2}\,(1+2\bar{n}_{a})\,\partial_{T}\bar{n}_{a}\right]^{2}C^{2}}{1-C^{2}},$ (6)
where $\partial_{T}\bar{n}_{a}=\hbar\omega_{a}\!\left[\bar{n}_{a}(\bar{n}_{a}+1)\right]/k_{B}T^{2}$. Eq. 6 establishes the fundamental precision bound of the CM thermometer [46, 40, 67, 17, 12].
Physically, the probe converts temperature-dependent photon-number fluctuations into measurable phase diffusion. In this CM model, the QFI increases with $\alpha^{2}$ but is simultaneously limited by visibility decay in the denominator. In the weak-dephasing regime ($\Gamma_{\phi}\ll 1$), one finds $\mathcal{F}_{C}\;\approx\;\alpha^{2}(\lambda\tau)^{2}\,\bar{n}_{a}\left(\hbar\omega_{a}/k_{B}T^{2}\right)^{2}.$ Thus, the coherent-probe scheme provides a tunable transduction “gain” via $\alpha$ (and $\tau$), but this gain is bounded in practice by visibility decay, available probe power, bandwidth, and backaction. We therefore interpret the improvement as a resource-assisted sensitivity, which is achieved by allocating probe photons to strengthen the temperature-to-phase conversion, rather than as an unlimited amplification.
Fig. 2(a) shows a simulated QFI, $\mathcal{F}_{C}$, heatmap, highlighting the optimal interaction time $\tau(T)$ shown as white dashed line for the maximized QFI at each temperature.
We define the temperature sensitivity as $\delta T_{\mathrm{min}}=1/\sqrt{\nu\mathcal{F}_{C}(T)},$ with $\nu$ the number of independent repetitions. Fig. 2(b) shows that the thermometer can achieve a sensitivity $\delta T\sim 60~\mu\mathrm{K}$ at $T=10~\mathrm{mK}$, with an optimal interaction time around $\tau\sim 25~\mu\mathrm{s}$ and measurement repetition $\nu=10^{4}$. Such timescales are readily supported by high-Q 3D cavities, which exhibit photon lifetimes up to millisecond range [49, 51, 36].
The calculated QFI as a function of $\tau$ for different cross-Kerr strengths is shown in Fig. 2(c), with fixed temperature $T$ = 10 mK. Since the qubit remains in its ground state during the sensing window $\tau$ and is only activated briefly afterward to map out the probe coherence (Appendix. G), this scheme provides a useful tuning knob to balance the cross-Kerr strength $\lambda$ and the interaction time $\tau$. A longer sensing window allows operation with a smaller $\lambda$, which in turn suppresses probe-induced backaction on the thermal absorber and permits a larger probe amplitude $\alpha$. Conversely, a stronger $\lambda$ enables faster sensing with shorter $\tau$, while maintaining moderate backaction by reducing $\alpha$ [50].
II.4 Strategy 2: Probe Phase-Shift (PS) Tracking